OpenStudy (anonymous):

After a devastating winter, when thousands of fish died, an environmental scientist has replenished the trout stock in a fishing pond. He started with 8,000 baby trout and has finished a count to find that, in 5 years, the population is estimated to be 24,000. Assuming an exponential growth pattern, what is the annual growth rate (rounded to the nearest tenth of a percent) of the new trout population? Hint: A(t) = A0(1+r)^t, where A(t) is the final amount, A0 is the initial amount, r is the growth rate expressed as a decimal, and t is time. A. 33.3% B. 2.5% C. 60.0% D. 24.6%

OpenStudy (xapproachesinfinity):

you just need to solve for r in that equation \[A(t)=A_0 (1+r)^t\] in five years we have \[A(5)=8000(1+r)^5\] \[24000=8000(1+r)^5 \]

OpenStudy (xapproachesinfinity):

can you solve for r?

OpenStudy (anonymous):

I think so

OpenStudy (anonymous):

\[r = -1 + \sqrt[5]{3}\]

OpenStudy (anonymous):

@xapproachesinfinity Is that right?

OpenStudy (xapproachesinfinity):

so what is the value?

OpenStudy (anonymous):

I'm not 100% sure. Could you maybe walk me through this? I don't really understand. If you have time at least

OpenStudy (xapproachesinfinity):

you already did half the work just use calculator

OpenStudy (anonymous):

Ok one sec

OpenStudy (anonymous):

I got 0.2457

OpenStudy (xapproachesinfinity):

you need to multiply by 100

OpenStudy (anonymous):

So 24.57

OpenStudy (xapproachesinfinity):

so?

OpenStudy (anonymous):

So the last choice?

OpenStudy (xapproachesinfinity):

yes

OpenStudy (anonymous):

Thank you for your help